2:25 Slewing to collimation target

not sending zernikes on DX. Seeing is 2.51"

2:28 Zernikes on DX do not look unreasonable so going to send them. Looking better with each iteration. Seeing is now showing 1.2-2" on SX

email sent to AOG with snap shot of WFS
SXvsDX WFS in poor seeing:
SXvsDX.png

With Good Seeing luci2 IQ was severely degraded. DX IQ with <1” was unusable, with LUCI images 4pixels on luci1.20220414.0071 and 10 pixels on luci2..0067.

DX IQ with good seeing (0.9" on teh DIMM, 0.65" on SX guider)
Screen Shot 2022-04-13 at 11.46.38 PM.png

2:37 OSCO on BD+70 622
  • OCAM magnitude 9.1
  • bin 1 900 Hz, modulation3, OCAM gain 500,
  • getting bite during optimize gain, 2.27" seeing probably to blame
  • We are into twilight so moving on to Jochen's seeing limited science

6:55 Sending ESM preset, AOref catalog magnitude is 12.6
  • OCAM magnitude 15.4
  • Safeskip during optimize gain.
  • Stopped and made sure pupils were well aligned by applying a little gain manually
7:02:00 Shell rip during optimize gain. Lots of safe skips from optimize gain, safeskip 12% still showed after failure.
  • fast_diag did not come up during adsc_start first time, ok second time.
  • Flat set with 503 modes

7:15 Problem with tip tilt request on psf left side while recovering the adsec. very odd. This persisted into the next preset. Josh cleared, but had to execute a pointing check afterwards cince star was lost. Collimation was also way out of whack.

7:24 Preset to ESM target with AOref catalog magnitude 12.6, clear conditions with variable seeing 0.9-1.3"
  • Seeing 0.93" on the DIMM, 0.6" on the LUCI1 GCS before starting
  • Bin4, 400Hz, modulation 3, OCAM gain 600
  • OCAM mag 15.3
  • Safeskip when optimizing gains during tiptilt. Optimized manually.
  • Gains 0.9, 0.5
  • TN 073507
  • Seeing on OCAM is not updating to match DIMM seeing so closed and reopened and it updated again.
8:01 Preset to TTM target Aoref catalog magnitude is 15.9
  • Target is 35degrees from the moon but is hard to discern from sky background. Failed center star. Resent target as a 17.5 which may be closer to its actualy magnitude (maybe closer to 18th) but it still did not pick target from background. Suggest this target be attempted on a darker night or with a further moon distance if at all.
8:10 Preset to TTM with AOref magnitude of 15.1
  • pointing check required.
  • Star is very faint again, possibly 3 mag difference. Also rerotator out of range so Josh is rotating 360
  • 8:21 sending preset with adjusted magnitude in script to 18
  • This is too faint to pull in. This AOref is not in UCAC5, only in nomad so likely a galaxy. It is visible on CCD47 but barely. Suggested i try to pull in with baysides but Jochen would like to just move on. PI knew the risks
8:26 Preset to TTM target with AOref catalog magnitude13.8
  • OCAM magnitude 15.2
  • bin4 400Hz OCAM gain 600, seeing 0.9-1.3" on the DIMM, clear
  • Optimize gain failed gain, this time on second round of TT, manually optimized: 0.80, 0.85, slope rms 0.5
  • TN 083744
9:12 Preset to TTM target with AOref catalog magnitude 12.79
  • OCAM magnitude 14.2
  • Bin2 400Hz, 600 OCAM gain
  • Gains 2.33, 8.70, 0.08 Slope RMS 0.67
  • TN 091948
  • Seeing 1.3-1.75" on the DIMM, clear
9:53 Preset to TTM target with AOref catalog magnitude 12.13
  • Center star sent the star off of ccd47
  • Resending preset. This time senter star worked fine. Really not sure what happened there...
  • OCAM magnitude 14.0
  • Bin2 400Hz Ocam Gain 600
  • Gains 2.36, 10.77, 0.08
  • TN 100603
  • Seeing 1.07" ont he DIMM, clear
10:40 AOref Catalog magnitude 12.3 (sent aas 12.6 to push bin2)
  • Y is railed bouncin betwee 100 and 92, rotator position is 389. Pupil differences are <0.1
  • OCAM magnitude 13.6
  • bin2, freq 400, OCAM gain 600
  • Seeing 1.05" on the DIMM, clear
  • Gains 4.17 9.00, 0.08
  • TN 105138 (camera len railed in Y
  • Offset failed to resume after pause. Naural guiding resumed. AO ref well within patrol field
    XStage:
    CopleyCtrl _171435 |ERR| 1577|2022-04-14 11:03:14.340575| MAIN > Error: Cannot write variable baysidex.L.STATUS: [AOVarException] Error writing variable (code -5001) Timeout error File RTDBvar.cpp line 496Variable: baysidex.L.STATUS (type INT_VARIABLE, len 1)
    CopleyCtrl _171435 |ERR| 1577|2022-04-14 11:03:14.340575| MAIN > .
    Ystage

    CopleyCtrl _171451 |ERR| 1965|2022-04-14 11:03:14.041789| MAIN > Error: Cannot write variable baysidey.L.STATUS: [AOVarException] Error writing variable (code -5001) Timeout error File RTDBvar.cpp line 496Variable: baysidey.L.STATUS (type INT_VARIABLE, len 1)
    CopleyCtrl _171451 |ERR| 1965|2022-04-14 11:03:14.041789| MAIN > .
    Zstage

    CopleyCtrl_171467 |ERR| 235|2022-04-14 11:03:14.799723| MAIN > Error: Cannot write variable baysidez.L.STATUS: [AOVarException] Error writing variable (code -5001) Timeout error File RTDBvar.cpp line 496Variable: baysidez.L.STATUS (type INT_VARIABLE, len 1)
    CopleyCtrl_171467 |ERR| 235|2022-04-14 11:03:14.799723| MAIN > .

  • Next offset sent and AOref star returned to FOV
  • Pupil difference is now 0.3 -0.4 in Y. Camera lens is still railed (and of course when I hit start on TN the pupil diffeence dropped to 0: TN 111603)
  • 11:28
    Offset put Xstage outside rage

    171435 |ERR| 1728|2022-04-14 11:23:10.235001| MAIN > Error: movement out of range
    CopleyCtrl_171435 |ERR| 1729|2022-04-14 11:23:10.235061| MAIN > Request: -108.327996 Limits: 0.000000 -108.000000
    CopleyCtrl_171435 |ERR| 1730|2022-04-14 11:23:10.235153| MAIN > Variable handler error: [AOException] Error in MoveTo (): movement out of range (code -5002) Value is out of range
    CopleyCtrl_171435 |ERR| 1731|2022-04-14 11:23:53.330250| MAIN > Error: movement out of range
    CopleyCtrl_171435 |ERR| 1732|2022-04-14 11:23:53.330327| MAIN > Request: -108.327996 Limits: 0.000000 -108.000000
    CopleyCtrl_171435 |ERR| 1733|2022-04-14 11:23:53.330411| MAIN > Variable handler error: [AOException] Error in MoveTo (): movement out of range (code -5002) Value is out of range
    CopleyCtrl_171435 |ERR| 1734|2022-04-14 11:24:36.461700| MAIN > Error: movement out of range
    CopleyCtrl_171435 |ERR| 1735|2022-04-14 11:24:36.461791| MAIN > Request: -108.327996 Limits: 0.000000 -108.000000
    CopleyCtrl_171435 |ERR| 1736|2022-04-14 11:24:36.461888| MAIN > Variable handler error: [AOException] Error in MoveTo (): movement out of range (code -5002) Value is out of range

11:37 Preset to TTM target with AO ref catalog mag 13.7
  • OCAM mag 14.2
  • DIMM seeing 0.99
  • bin2, 400Hz, OCAM gain 600
  • Gains 3.74, 8.7, 0.08
  • TN 114128
12:20 Set safe state, ocame warming and shutting down for the night.

-- JenniferPower - 14 Apr 2022
Topic revision: r2 - 14 Apr 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback